"\int" Cosec6(2x) dx


1
Expert's answer
2021-01-01T12:42:05-0500

Let

"I=\int cosec^6(2x)dx"

Let "y=2x\Rightarrow dy=2dx"


Now "I=\int cosec^6(2x)dx=\frac{1}{2}\int cosec^6(y)dy"

"\therefore 2I\\= \int cosec^6ydy\\
=\int cosec^4y\times cosec^2ydy\\
=\int cosec^4y\times(1+cot^2y)dy\\
=\int (cosec^4y)dy+\int cosec^4y\times cot^2ydy\\
=\int (cosec^2y\times(1+ cot^2y))dy+\int cosec^2y\times cot^2y\times (1+cot^2y)dy\\
=\int cosec^2ydy+\int cosec^2ycot^2ydy+\int cosec^2ycot^2ydy+\int cot^4ycosec^2ydy\\"

Let

"coty=z\\\Rightarrow -cosec^2ydy=dz\\\Rightarrow cosec^2ydy=-dz"


Now replacing "coty \;by\; z\;and \;cosec^2ydy\;by \;(-dz),"

"2I\\
=-\int dz-2\int z^2dz-\int z^4dz\\
=-z- \frac{2z^3}{3}-\frac{z^5}{5}+c'"

[ c' is the integrating constant ]

"=-cot(2x)-\frac{2}{3}cot^3(2x)-\frac{1}{5}cot^5(2x)+c'"

[ putting "z=coty" and, "y=2x" ]

"\Rightarrow I=-\frac{1}{2}cot(2x)-\frac{2}{6}cot^3(2x)-\frac{1}{10}cot^5(2x)+c"

"c=c'\/2]"


So,


"\boxed{I=\int cosec^6(2x)=-\frac{1}{2}cot(2x)-\frac{2}{6}cot^3(2x)-\frac{1}{10}cot^5(2x)+c}"

where "c" is a Integrating Constant.